Lowell Company makes and sells artistic frames for pictures. The controller is responsible for preparing the master budget and has accumulated the following information for 2020.

irx1TGM7cU2NivfJmeXQzLNx4y1sWzy15sKeNdGc

Lowell has a labor contract that calls for a wage increase to $10 per hour on April 1. New labor-saving machinery has been installed and will be fully operational by March 1.

Lowell expects to begin the year with 17,580 frames on hand and has a policy of carrying an end-of-month inventory of 100% of the following month’s sales, plus 60% of the second following month’s sales.

Prepare a direct labor budget for Lowell Company by month and for the first quarter of the year. The direct labor budget should include direct labor hours.

Quhad6rQL0WmeTlMyvgAAGmV8AQA0yPgCAGiQ8QU
Answer #1
Direct Labor Budget
For the Quarter Ending March 31, 2020
January February March Total
Production in Units 9760 8160 8120
Direct Labor Hours Per Unit 2.2 2.2 1.5
Total labor Hours Needed 21472 17952 12180
Rate Per Hour 9 9 9
Total Direct Labor 193248 161568 109620 464436
Workings:
Production Budget
January February March
Sales in Units 10500 11800 8400
Add: Desired Ending Inventory 16840 13200 12920
Total Needs 27340 25000 21320
Less: Beginning Inventory 17580 16840 13200
Required Production Units 9760 8160 8120
Desired Ending Inventory
January 16840 =11800+(8400*60%)
February 13200 =8400+(8000*60%)
March 12920 =8000+(8200*60%)